OpenStudy (anonymous):

Help me With thiss PLEASE!!! lim 1-2x^2-2cosx+cos^2x/x^1/2 approaches x -> 0

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-2x^2-2\cos(x)+\cos^2(x)}{x^\frac{1}{2}}\]

OpenStudy (freckles):

this might be helpful \[\cos^2(x)-2\cos(x)+1=(\cos(x)-1)^2 \]

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(\cos(x)-1)^2-2x^2}{x^\frac{1}{2}} \\ \lim_{x \rightarrow 0}\frac{x^\frac{3}{2}}{x^\frac{3}{2}}\frac{(\cos(x)-1)^2-2x^2}{x^\frac{1}{2}} \\ \lim_{x \rightarrow 0}\frac{ x^\frac{3}{2}(\cos(x)-1)^2-x^\frac{3}{2}2x^2}{x^2} \\ \lim_{x \rightarrow 0} x^\frac{3}{2} \frac{(\cos(x)-1)^2}{x^2}-\lim_{x \rightarrow 0}\frac{x^\frac{3}{2} 2x^2}{x^2} \\ \lim_{x \rightarrow 0}x^\frac{3}{2} (\lim_{x \rightarrow 0} \frac{\cos(x)-1}{x})^2-\lim_{x \rightarrow 0}2x^\frac{3}{2}\]

OpenStudy (freckles):

and I kinda think they mean 0^+ since x^(1/2) doesn't actually exist for the real numbers when x<0

OpenStudy (xapproachesinfinity):

must be x>0

OpenStudy (anonymous):

In my answer sheet appear -2 D: Im confused S:

OpenStudy (freckles):

I wrote down the function you wanted right?

OpenStudy (freckles):

I don't see how they get -2

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(\cos(x)-1)^2-2x^2}{x^\frac{1}{2}} \\ \lim_{x \rightarrow 0}\frac{x^\frac{3}{2}}{x^\frac{3}{2}}\frac{(\cos(x)-1)^2-2x^2}{x^\frac{1}{2}} \\ \lim_{x \rightarrow 0}\frac{ x^\frac{3}{2}(\cos(x)-1)^2-x^\frac{3}{2}2x^2}{x^2} \\ \lim_{x \rightarrow 0} x^\frac{3}{2} \frac{(\cos(x)-1)^2}{x^2}-\lim_{x \rightarrow 0}\frac{x^\frac{3}{2} 2x^2}{x^2} \\ \lim_{x \rightarrow 0}x^\frac{3}{2} (\lim_{x \rightarrow 0} \frac{\cos(x)-1}{x})^2-\lim_{x \rightarrow 0}2x^\frac{3}{2}=0(0)^2-2(0) \\ =0-0=0\]

OpenStudy (freckles):

assuming we are looking at the right limit of course

OpenStudy (anonymous):

wait damn it I made a mistake the denominator is x^2 !! not x^1/2 D:

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-2x^2-2\cos(x)+\cos^2(x)}{x^2} \\ \lim_{x \rightarrow 0}\frac{(\cos(x)-1)^2-2x^2}{x^2}\] separate the fraction it isn't too much work after that

OpenStudy (anonymous):

hahahaha Oh what a wonderful math night S: I almost got insane

OpenStudy (freckles):

math can make one insane

OpenStudy (freckles):

just check out Einstein's hair

OpenStudy (freckles):

anyways I must go you have fun

OpenStudy (anonymous):

THANKSS :D jajaja